# Automatic Tracing in Task-Based Runtime Systems

### v1

@@ 1. Introduction
In order to improve programmer productivity and to enable the tracing of modular high-level programs, implicitly parallel task-based systems should automatically identify repeated sequences of operations, memoize their analysis results and cheaply replay the analysis as needed. We call this the problem of automatic trace identification, which is similar to Just-In-Time (JIT) compilation in the context of dynamic language implementations (Hölzle and Ungar, 1996; Gal et al., 2009; Paleczny et al., 2001). JIT compilers for dynamic languages interpret bytecode during program startup, and compile bytecode to native instructions as repeatedly invoked program fragments become hot. Following this architecture, implicitly parallel task-based runtimes should interpret issued operations with a dynamic dependence analysis, and switch to an analysis-free compiled execution once repeated sequences of operations are encountered.
@@ 2. Motivating Example
We now show an example of high-level implicitly parallel code where it is difficult for a programmer place tracing annotations. As part of developing the example, we provide necessary background on the Legion (Bauer et al., 2012) runtime system.
Figure 1(a) performs Jacobi iteration using cuNumeric (Bauer and Garland, 2019), a distributed drop-in replacement for NumPy. cuNumeric distributes NumPy through a dynamic translation to Legion. cuNumeric implements NumPy operations by issuing one or more Legion tasks, which are designated functions registered with the runtime system. Each NumPy array is mapped to a Legion region, which is a multi-dimensional array tracked by Legion. Each task takes a list of regions as arguments. The stream of tasks launched by the main loop of the cuNumeric program is in Figure 1(b). For each task, the first two arguments denote the inputs, while the third argument is the output. Legion extracts parallelism from the issued stream of tasks by analyzing the data dependencies between tasks and the usage of their region arguments (Bauer et al., 2023).
1import cunumeric as np
@@
Figure 1. A cuNumeric (Bauer and Garland, 2019) program and the stream of tasks it issues at runtime. An intuitive trace around the main loop does not correspond to a repeated program fragment.
@@
A natural attempt to trace the program in Figure 1(a) would place the tbegin and tend around the body of the main for loop. However, this annotation results in an invalid trace (raising a runtime error), for a subtle reason that requires knowledge of the internals of cuNumeric. The problem with this natural annotation is the loop-carried use of the Python variable x, which is bound to different cuNumeric arrays (regions) at different points of execution. Upon entering loop iteration $i$, x is bound to a region arbitrarily named x1, which is used as an argument for the first dot operation. As execution proceeds, cuNumeric allocates a new region x2 for the result of the division with d, and binds the Python variable x to the region x2. Therefore, the next iteration $i+1$ issues a dot on x2, causing iteration $i+1$ to issue a different sequence of tasks than iteration $i$! This program illustrates a real-world case where abstraction and composition make it difficult to apply the low-level tracing technique.
To correctly trace the program in Figure 1(a), a programmer must either add trace annotations around every two iterations of the main loop, or use two different trace ID’s for each different iteration’s repetition pattern. This steady state of groups of two iterations is achieved because when x is assigned, the region it refers to can be collected and immediately reused by cuNumeric. Relying on this steady state is brittle, as the addition of more operations in the main loop or a change in cuNumeric’s region allocation policy could perturb the way in which the necessary steady state for tracing is achieved. Instead, Apophenia dynamically analyzes the stream of tasks and automatically discovers what fragments of the application should be traced, removing this concern from the programmer.
@@ 4. Trace Identification
Dynamically finding good traces requires processing information about the tasks seen so far, and then using that information to record and replay traces in the future. An overview of Apophenia’s dynamic analysis procedure is sketched in Algorithm 1. Apophenia has two components that correspond to the to the targets of the optimization problem in Section 3. The trace finder constructs the candidate set of traces $T$ by accumulating the tasks issued by the application into a buffer, and asynchronously mining the buffer to find candidate traces. The trace replayer then constructs the matching function $f$ by ingesting the candidate traces into a trie, and identifying candidate traces in the application stream by maintaining pointers into the trie that represent potential matches. A concrete example of how Apophenia identifies a trace in an application is shown in Figure 3. We now describe each of these components in detail.
@@ 4.2. Finding Traces With High Coverage
Apophenia’s trace finder records tasks as they are issued by the application into a buffer (we describe a refinement to this scheme in Section 4.4). Once the buffer fills up, Apophenia launches an asynchronous analysis of the buffer to find a set of traces within the buffer that maximize the coverage of the buffer. We discuss previous ideas that are related to this goal, and then describe the solution used in Apophenia. [^2] We discuss more related works in Section 7.
@@
Within the programming languages community, recent work by Sisco et al. (Sisco et al., 2023) used a technique called tandem repeat analysis (Stoye and Gusfield, 2002) to find loops in the netlists that result from compiling hardware description languages. A tandem repeat is a sub-string $\alpha$ that repeats contiguously within a larger string $S$, such that $\alpha^{k}$ is a sub-string of $S$, for some $k$. Despite the success that Sisco et al. had using tandem repeat analysis, we found that even simple real world cuNumeric programs did not contain enough tandem repeats for the analysis to reliably identify a trace set with high coverage. The reason is that while these real-world programs tended to have repetitive main loops, there would often be irregularly appearing computations such as convergence checks or statistics calculations that occur infrequently between loop iterations. As such, the strings that represented these programs would not contain tandem repeats, but instead repeated sub-strings separated by other tokens.
@@
In this work, we design a repeat finding algorithm that is directly aware of the optimization problem in Section 3 and runs in $O(n\log(n))$, where $n$ is the size of the token history buffer. At a high level, our algorithm makes a pass through a suffix array constructed from the input buffer to collect a set of candidate repeats. It then greedily selects the largest repeated sub-strings that do not overlap with any previously chosen sub-strings. Psuedocode for our algorithm is in Algorithm 2, which takes a string $S$ and returns a set of sub-strings that achieve high coverage of $S$. We assume that the reader is knowledgeable about suffix arrays and their structural properties. However, understanding the algorithm in Algorithm 2 is not required to understand its usage in Apophenia, as discussed in Section 4.3 and Section 4.4.
@@ 24                   $R\leftarrow R+[S[s:s+l]]$
Our algorithm aims to find good solutions to the optimization problem in Section 3 by identifying long repeated sub-strings and selecting as many as possible that do not overlap with each other. We trade off between an optimal solution to the optimization problem to instead find good solutions and maintain a lower asymptotic runtime. There are two such heuristics in our algorithm. First, when adjacent suffix array entries have a repetition, we consider only the maximal length repetition instead of all sub-strings of the repetition. Second, when we select which candidates to keep, we greedily choose the largest candidates instead of performing a bin-packing computation. While we do not provide theoretical guarantees on the optimality of our algorithm, we show in Section 6 that Apophenia using our algorithm is able to identify good traces in complex, real-world applications.
@@ 4.3. Recognizing and Replaying Candidate Traces
Apophenia’s trace replayer uses Algorithm 2 to find candidate traces from the application’s history of tasks. In this section, we discuss how Apophenia’s trace replayer identifies and selects these candidate traces from the task stream to record and replay. Our design of the trace replayer has two major goals. First, the per-task overhead must be low, as it is imperative for performance for the application to issue as many tasks into the runtime as possible so that the runtime can either replay traces or perform dependence analysis ahead of execution. Slowing down the task launch rate would result in exposed latency from various sources in the runtime. Second, Apophenia must balance exploration and exploitation when selecting traces. As more information about the application is gained, Apophenia should switch to better traces as it finds them. However, Apophenia should not leave a steady state until it is confident that performance can be improved, as recording new traces has a cost.
@@ 5.1. Distributing the Analysis
Apophenia’s analysis as presented in Section 4 is sequential, processing tasks as they are issued by the application. In a distributed setting, Apophenia leverages Legion’s dynamic control replication (Bauer et al., 2021) to act as a sequential analysis, except for one component, which we discuss next. With control replication, the application executes on each node and Legion shards the dependence analysis and execution across nodes. The main restriction of control replication is that the application must issue the same sequence of tasks on every node. We implement Apophenia as a layer between the application and Legion, so Apophenia inherits the control replication requirements of the application. In particular, each node must agree on which traces to replay and when during program execution to record and replay the traces.
The only source of non-determinism in Apophenia that may result in divergent decisions between nodes is the asynchronous processing of token buffers described in Section 4.2. The asynchronous analysis may complete earlier on one node than another, resulting in that node replaying a trace before another node has identified that trace as a candidate. However, making the analysis synchronous would result in stalling the application until analyses complete. We resolve this tension by having each node agree on a count of processed operations to issue before ingesting the results of an asynchronous analysis. If any node had to wait on an asynchronous analysis to complete, all nodes increase their count of operations to wait on for the next analysis. This strategy reaches a steady state where analysis results are ingested in a deterministic manner without stalling the application.
@@ 6. Evaluation
Figure 6. Weak scaling on previously traced Legion applications, where Apophenia (“auto” in the figure) performs competitively.
@@
Figure 7. Weak scaling on cuNumeric applications, where Apophenia (“auto” in the figure) outperforms the untraced version.
@@
We evaluated Apophenia on the Eos and Perlmutter supercomputers. Each node of Eos is an NVIDIA DGX H100, containing 8 H100 GPUs with 80 GB of memory and a 112 core Intel Xeon Platinum. Each node of Perlmutter contains 4 NVIDIA A100 GPUs with 40 GB of memory and a 64 core AMD EPYC 7763. Nodes of Eos are connected with an Infiniband interconnect, while Perlmutter uses a Slingshot interconnect. We compile Legion on Eos with the UCX networking module, and use the GASNet-EX (Bonachea and Hargrove, 2018) networking module on Perlmutter. We do not execution each application on both Perlmutter and Eos due to differences between the local environments on each machine. In our experiments, we evaluate the relative performance differences between traced and untraced programs, and comparisons between machines are not significant.
@@ 6.1. Weak Scaling
S3D (Treichler et al., 2017) is a production combustion chemistry simulation code that has been developed over the course of many years by different scientists and engineers. The Legion port of S3D implements the right-hand-side function of the Runge-Kutta scheme, and interoperates with the legacy Fortran+MPI driver of the simulation. The integration between Legion and the legacy Fortran+MPI code leads to various constraints that the manual trace annotations interact with. For example, during the first 10 iterations, a hand-off between Legion and Fortran+MPI must occur every iteration, while after the first 10 iterations a hand-off is required only every 10 iterations. While not unmanageable, these interactions have led to relatively complicated logic to manually trace the main loop. We scale S3D on Perlmutter, and compare the performance of Apophenia to manually traced and untraced versions of S3D. The results are shown in Figure 6(a). Even on a single node, tracing has a noticeable performance impact on the smaller problem sizes and affects the scalability of S3D. Apophenia achieves within 0.92x–1.03x of the performance of the manually traced version, and between 0.98x–1.82x speedups over the untraced version.
@@
CFD is a cuNumeric application that solves the Navier-Stokes equations for 2D channel flow (Barba and Forsyth, 2019). Unlike S3D and HTR, there is not a manually traced version of CFD, due to the difficulties around composition discussed in Section 2. Developing a manually traced implementation of CFD would require either rewriting the application to remove any dynamic region allocation, or manual examination of allocator logs to find the number of iterations in the steady state. As a result, we compare CFD with Apophenia to the standard untraced version on different problem sizes, which is the performance that cuNumeric users are able to achieve today.
@@
TorchSWE is a cuNumeric port of the MPI-based TorchSWE (Chuang, 2021) shallow-water equation solver, and is the largest cuNumeric application developed so far. Similarly to CFD, there is no manually traced version to compare to. However, unlike CFD, performing a rewrite of TorchSWE to enable manual tracing would be difficult, as TorchSWE contains an order of magnitude more lines of code. Weak scaling results for TorchSWE on Eos are shown in Figure 7(b).
These results demonstrate that there does not exist a problem size for TorchSWE on Eos that can hide Legion’s runtime overhead without tracing. Even the large problem size, which nearly reaches the GPU’s memory capacity, exposes Legion runtime overhead at 8 GPUs. The reason for this is that TorchSWE maintains a large number of fields for each simulated point, and issues different array operations on each field. The amount of data needed for each element in the simulation does not allow the task granularity to be easily increased, as each new element added increases the memory footprint more than it increases the average task granularity. For such applications, leveraging tracing is a requirement, and Apophenia enables complex applications like TorchSWE to do so automatically. TorchSWE itself contains enough task parallelism to hide communication latencies, but needs tracing to first lower runtime overhead. With Apophenia, we are able to achieve between 0.91x–2.82x speedup on TorchSWE, achieving nearly perfect scalability on 64 GPUs.

### v2

@@ Abstract.
Dynamic Analysis; Runtime Systems; Tracing
journalyear: 2025copyright: ccconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 30-April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’25), March 30-April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3669940.3707237isbn: 979-8-4007-0698-1/25/03ccs: Computing methodologies Distributed programming languages
@@ 1. Introduction
In order to improve programmer productivity and to enable the tracing of modular high-level programs, implicitly parallel task-based systems should automatically identify repeated sequences of operations, memoize their analysis results and cheaply replay the analysis as needed. We call this the problem of automatic trace identification, which is similar to Just-In-Time (JIT) compilation in the context of dynamic language implementations (Hölzle and Ungar, 1996; Gal et al., 2009a; Paleczny et al., 2001). JIT compilers for dynamic languages interpret bytecode during program startup, and compile bytecode to native instructions as repeatedly invoked program fragments become hot. Following this architecture, implicitly parallel task-based runtimes should interpret issued operations with a dynamic dependence analysis, and switch to an analysis-free compiled execution once repeated sequences of operations are encountered.
@@ 2. Motivating Example
1import cupynumeric as np
@@
Figure 1. A cuPyNumeric (Bauer and Garland, 2019) program and the stream of tasks it issues at runtime. An intuitive trace around the main loop does not correspond to a repeated program fragment.
We now show an example of high-level implicitly parallel code where it is difficult for a programmer to place tracing annotations. As part of developing the example, we provide necessary background on the Legion (Bauer et al., 2012) runtime system.
Figure 1(a) performs Jacobi iteration using cuPyNumeric (Bauer and Garland, 2019), a distributed drop-in replacement for NumPy. cuPyNumeric distributes NumPy through a dynamic translation to Legion. cuPyNumeric implements NumPy operations by issuing one or more Legion tasks, which are designated functions registered with the runtime system. Each NumPy array is mapped to a Legion region, which is a multi-dimensional array tracked by Legion. Each task takes a list of regions as arguments. The stream of tasks launched by the main loop of the cuPyNumeric program is in Figure 1(b). For each task, the first two arguments denote the inputs, while the third argument is the output. Legion extracts parallelism from the issued stream of tasks by analyzing the data dependencies between tasks and the usage of their region arguments (Bauer et al., 2023).
@@
A natural attempt to trace the program in Figure 1(a) would place the tbegin and tend around the body of the main for loop. However, this annotation results in an invalid trace , for a subtle reason that requires knowledge of the internals of cuPyNumeric. The problem with this natural annotation is the loop-carried use of the Python variable x, which is bound to different cuPyNumeric arrays (regions) at different points of execution. Upon entering loop iteration $i$, x is bound to a region arbitrarily named x1, which is used as an argument for the first dot operation. As execution proceeds, cuPyNumeric allocates a new region x2 for the result of the division with d, and binds the Python variable x to the region x2. Therefore, the next iteration $i+1$ issues a dot on x2, causing iteration $i+1$ to issue a different sequence of tasks than iteration $i$! Issuing a different sequence of tasks with the same trace id is a violation of the conditions to use tracing, and the runtime system may either raise an error or fall back to the expensive dependence analysis. This program illustrates a real-world case where abstraction and composition make it difficult to apply the low-level tracing technique.
To correctly trace the program in Figure 1(a), a programmer must either add trace annotations around every two iterations of the main loop, or use two different trace ID’s for each different iteration’s repetition pattern. This steady state of groups of two iterations is achieved because when x is assigned, the region it refers to can be collected and immediately reused by cuPyNumeric. Relying on this steady state is brittle, as the addition of more operations in the main loop or a change in cuPyNumeric’s region allocation policy could perturb the way in which the necessary steady state for tracing is achieved. Instead, Apophenia dynamically analyzes the stream of tasks and automatically discovers what fragments of the application should be traced, removing this concern from the programmer.
@@ 4. Trace Identification
Dynamically finding good traces requires processing information about the tasks seen so far, and then using that information to record and replay traces in the future. An overview of Apophenia’s dynamic analysis procedure is sketched in Algorithm 1. Apophenia has two components that correspond to the targets of the optimization problem in Section 3. The trace finder constructs the candidate set of traces $T$ by accumulating the tasks issued by the application into a buffer, and asynchronously mining the buffer to find candidate traces. The trace replayer then constructs the matching function $f$ by ingesting the candidate traces into a trie, and identifying candidate traces in the application stream by maintaining pointers into the trie that represent potential matches. Apophenia intercepts calls to target runtime’s ExecuteTask function, and forwards a potentially different set of tasks and trace markers to the runtime. A concrete example of how Apophenia identifies a trace in an application is shown in Figure 3. We now describe each of these components in detail.
@@ 4.2. Finding Traces With High Coverage
Apophenia’s trace finder records tasks as they are issued by the application into a buffer (we describe a refinement to this scheme in Section 4.4). Once the buffer fills up, Apophenia launches an asynchronous analysis of the buffer to find a set of traces within the buffer that maximize the coverage of the buffer. We discuss previous ideas that are related to this goal, and then describe the solution used in Apophenia. [^2] We discuss more related work in Section 7.
@@
Within the programming languages community, recent work by Sisco et al. (Sisco et al., 2023) used a technique called tandem repeat analysis (Stoye and Gusfield, 2002) to find loops in the netlists that result from compiling hardware description languages. A tandem repeat is a sub-string $\alpha$ that repeats contiguously within a larger string $S$, such that $\alpha^{k}$ is a sub-string of $S$, for some $k$. Despite the success that Sisco et al. had using tandem repeat analysis, we found that even simple real world cuPyNumeric programs did not contain enough tandem repeats for the analysis to reliably identify a trace set with high coverage. The reason is that while these real-world programs tended to have repetitive main loops, there would often be irregularly appearing computations such as convergence checks or statistics calculations that occur infrequently between loop iterations. As such, the strings that represented these programs would not contain tandem repeats, but instead repeated sub-strings separated by other tokens.
@@
In this work, we design a repeat finding algorithm that is directly aware of the optimization problem in Section 3 and runs in $O(n\log(n))$, where $n$ is the size of the token history buffer. At a high level, our algorithm makes a pass through a suffix array constructed from the input buffer to collect a set of candidate repeats. It then greedily selects the largest repeated sub-strings that do not overlap with any previously chosen sub-strings. Psuedocode for our algorithm is in Algorithm 2333We also present a standalone implementation of the algorithm available at https://github.com/david-broman/matching-substrings., which takes a string $S$ and returns a set of sub-strings that achieve high coverage of $S$. We assume that the reader is knowledgeable about suffix arrays and their structural properties. However, understanding the algorithm in Algorithm 2 is not required to understand its usage in Apophenia, as discussed in Section 4.3 and Section 4.4.
@@ 24                   $R\leftarrow R+[S[s:s+l]]$
Our algorithm aims to find good solutions to the optimization problem in Section 3 by identifying long repeated sub-strings and selecting as many as possible that do not overlap with each other. We trade off between an optimal solution to the optimization problem to instead find good solutions and maintain a lower asymptotic runtime. There are two such heuristics in our algorithm. First, when adjacent suffix array entries have a repetition, we consider only the maximal length repetition instead of all sub-strings of the repetition. Second, when we select which candidates to keep, we greedily choose the largest candidates instead of performing a bin-packing computation. Our algorithm is guaranteed to find the longest repeated sub-string, but due to the second heuristic, we cannot provide theoretical guarantees about the other chosen sub-strings. We show in Section 6 that Apophenia using our algorithm is able to identify good traces in complex, real-world applications.
@@ 4.3. Recognizing and Replaying Candidate Traces
Apophenia’s trace replayer uses Algorithm 2 to find candidate traces from the application’s history of tasks. In this section, we discuss how Apophenia’s trace replayer identifies and selects these candidate traces from the task stream to record and replay. Our design of the trace replayer has two major goals. First, the per-task overhead must be low, as it is imperative for performance for the application to issue as many tasks into the runtime as possible so that the runtime can either replay traces or perform dependence analysis ahead of execution. Slowing down the task launch rate would result in exposed latency from various sources in the runtime. Second, Apophenia must balance exploration and exploitation when selecting traces. As more information about the application is gained, Apophenia should switch to better traces as it finds them. However, Apophenia should not leave a steady state of replaying a particular trace until it is confident that performance can be improved, as memoization of the dependence analysis for new traces has a cost.
@@ 5.1. Distributing the Analysis
Apophenia’s analysis as presented in Section 4 is sequential, processing tasks as they are issued by the application. In a distributed setting, Apophenia leverages Legion’s dynamic control replication (Bauer et al., 2021) to act as a sequential analysis, except for one component, which we discuss next. With control replication, the application executes on each node and Legion shards the dependence analysis and execution across nodes. The main restriction of control replication is that the application must issue the same sequence of tasks on every node. We implement Apophenia as a layer between the application and Legion, meaning that Apophenia intercepts calls into the Legion runtime from the application and forwards a (possibly different) set of calls into Legion. As such, Apophenia inherits the control replication requirements of the application. In particular, each node must agree on which traces to replay and when during program execution to record and replay the traces.
The only source of non-determinism in Apophenia that may result in divergent decisions between nodes is the asynchronous processing of token buffers described in Section 4.2. An instance of Apophenia exists on each node of the target machine, and each instance maintains a local history buffer of tasks to run asynchronous analyses on. The asynchronous analysis may complete earlier on one node than another, resulting in that node replaying a trace before another node has identified that trace as a candidate. However, making the analysis synchronous would result in stalling the application until analyses complete. We resolve this tension by having each node agree on a count of processed operations to issue before ingesting the results of an asynchronous analysis. If any node had to wait on an asynchronous analysis to complete, all nodes increase their count of operations to wait on for the next analysis. This strategy reaches a steady state where analysis results are ingested in a deterministic manner without stalling the application.
@@ 6. Evaluation
Figure 6. Weak scaling on previously traced Legion applications, where Apophenia (“auto”) performs competitively.
@@
Figure 7. Weak scaling on cuPyNumeric applications, where Apophenia (“auto”) outperforms the untraced version.
@@
We evaluated Apophenia on the Eos and Perlmutter supercomputers. Each node of Eos is an NVIDIA DGX H100, containing 8 H100 GPUs with 80 GB of memory and a 112 core Intel Xeon Platinum. Each node of Perlmutter contains 4 NVIDIA A100 GPUs with 40 GB of memory and a 64 core AMD EPYC 7763. Nodes of Eos are connected with an Infiniband interconnect, while Perlmutter uses a Slingshot interconnect. We compile Legion on Eos with the UCX networking module, and use the GASNet-EX (Bonachea and Hargrove, 2018) networking module on Perlmutter. We do not execute each application on both Perlmutter and Eos due to differences between the local environments on each machine. In our experiments, we evaluate the relative performance differences between traced and untraced programs, and comparisons between machines are not significant.
@@ 6.1. Weak Scaling
S3D (Treichler et al., 2017) is a production combustion chemistry simulation code that has been developed over the course of many years by different scientists and engineers. The Legion port of S3D implements the right-hand-side function of the Runge-Kutta scheme, and interoperates with the legacy Fortran+MPI driver of the simulation. The integration between Legion and the legacy Fortran+MPI code leads to various constraints that the manual trace annotations interact with. For example, during the first 10 iterations, a hand-off between Legion and Fortran+MPI must occur every iteration, while after the first 10 iterations a hand-off is required only every 10 iterations. While not unmanageable, these interactions have led to relatively complicated logic to manually trace the main loop. We scale S3D on Perlmutter, and compare the performance of Apophenia to manually traced and untraced versions of S3D. The results are shown in Figure 6(a). Even on a single node, tracing has a noticeable performance impact on the smaller problem sizes and affects the scalability of S3D. Apophenia achieves within 0.92x–1.03x of the performance of the manually traced version, and between 0.98x–1.82x speedups over the untraced version. Manual annotations can slightly outperform Apophenia by leveraging application knowledge to select traces that have lower replay overhead.
@@
CFD is a cuPyNumeric application that solves the Navier-Stokes equations for 2D channel flow (Barba and Forsyth, 2019). Unlike S3D and HTR, there is not a manually traced version of CFD, due to the difficulties around composition discussed in Section 2. Developing a manually traced implementation of CFD would require either rewriting the application to remove any dynamic region allocation, or manual examination of allocator logs to find the number of iterations in the steady state. As a result, we compare CFD with Apophenia to the standard untraced version on different problem sizes, which is the performance that cuPyNumeric users are able to achieve today.
@@
TorchSWE is a cuPyNumeric port of the MPI-based TorchSWE (Chuang, 2021) shallow-water equation solver, and is the largest cuPyNumeric application developed so far. Similarly to CFD, there is no manually traced version to compare to. However, unlike CFD, performing a rewrite of TorchSWE to enable manual tracing would be difficult, as TorchSWE contains an order of magnitude more lines of code. Weak scaling results for TorchSWE on Eos are shown in Figure 7(b), which show that TorchSWE is significantly impacted by Legion runtime overhead without tracing.
These results demonstrate that there does not exist a problem size for TorchSWE on Eos that can hide Legion’s runtime overhead without tracing. Even the large problem size, which nearly reaches the GPU’s memory capacity, exposes Legion runtime overhead at 8 GPUs. The reason for this is that TorchSWE maintains a large number of fields for each simulated point, and issues different array operations on each field. The amount of data needed for each element in the simulation does not allow the task granularity to be easily increased to the untraced Legion minimum of 1̃ms per task, as each new element added increases the memory footprint more than it increases the average task granularity. For such applications, leveraging tracing is a requirement, and Apophenia enables complex applications like TorchSWE to do so automatically. TorchSWE itself contains enough task parallelism to hide communication latencies, but needs tracing to first lower runtime overhead. With Apophenia, we are able to achieve between 0.91x–2.82x speedup on TorchSWE, achieving nearly perfect scalability on 64 GPUs.